OpenStudy (anonymous):

The perimeter of a rectangular garden is 40 meters. The length of the garden is 1 meter less than twice its width. What is the length of the garden?

OpenStudy (anonymous):

Hint: l=2w-1

OpenStudy (anonymous):

Now solve

OpenStudy (anonymous):

40=2(2w-1)+2w 40=4w-2+2w 40=6w-2 40+2=6w 42/6=w w=7 l=2w-1 l=2(7)-1 l=14-1 l=13